# From Bugs to Benchmarks: A Comprehensive Survey of Software Defect Datasets

### v1

@@ 3.3. Programming Languages of Defective Software
Figure 5. Breakdown of programming languages of defects in each application domain.
@@
Figure 5 shows the distribution of the most popular programming languages across different application domains. Python and C/C++ dominate machine learning-related defect datasets, and Java and Kotlin cover the major datasets for mobile applications. Compilers and runtime engines have a variety of programming languages, including C/C++, Rust, Python, JavaScript, and Go. Most blockchain-related defects are from Solidity smart contracts, except for one dataset that collects defects from blockchain clients (Kim et al., 2022), which are implemented in Java, Go, C#, and Rust. In other domains, such as autonomous systems, quantum computing, scientific computing, and computational notebooks, C/C++ and Python are commonly used, whereas datasets for cloud computing, databases, and the internet of things involve Java, Go, C/C++, Python, and JavaScript.
Figure 6. Breakdown of types of defects for each programming language.
@@
The types of defects are not equally distributed across programming languages, but instead some languages attract datasets of specific defect types. Figure 6 shows the distribution of software defect datasets across different programming languages and defect types. Java, C/C++, Python, JavaScript, and Go are common across all defect types, while C#, Rust, Solidity, and PHP cover functional, security, and performance defects only. Kotlin is the programming language with the least coverage of defect types in surveyed datasets, addressing only functional and security defects.
4. Dataset Construction
@@ 4.1. Source of Defects
The source of a defect refers to the origin or platform from which the defect data is collected. As listed in Table 5, there are diverse sources for dataset creators to mine defects from, such as issue reports, version control systems, vulnerability databases, developer Q&A forums, and others. Most of these sources are closely tied to the software development process, as illustrated in Figure 7, especially in software implementation, testing, deployment, and maintenance stages. Other sources, such as coding assignments and defect synthesis, offer a more controlled environment for collecting or creating defects. Among these sources, issue reports and version control systems are the most popular, accounting for 57 (57/132) and 40 (40/132) software defect datasets, respectively. Coding assignments, CI/CD, and defect synthesis are less frequently used, with only 6 (6/132) datasets utilizing each. The rest of this section provides a detailed discussion of the primary sources of defects in software defect datasets, highlighting their unique characteristics.
@@
Issue Reports (Timperley et al., 2024; Andrade et al., 2024; Liu et al., 2024b; Zhang et al., 2024a; Cui et al., 2024; Yu et al., 2024; Xu et al., 2024; Waseem et al., 2024; Das et al., 2024; Yu and Wang, 2024; Jimenez et al., 2024; Zhao et al., 2023a; Jiang et al., 2023; Applis and Panichella, 2023; Bhuiyan et al., 2023; Guan et al., 2023; Xiong et al., 2023; Pachouly et al., 2022; Liang et al., 2022; Keshavarz and Nagappan, 2022; Liu et al., 2022; Wendland et al., 2021; Kim et al., 2021b; Paul et al., 2021; Du et al., 2021; Makhshari and Mesbah, 2021; Ferenc et al., 2020; Wu et al., 2020; Garcia et al., 2020; Vieira et al., 2019; Riganelli et al., 2019; Ponta et al., 2019; Saha et al., 2018; Zhang et al., 2018; Ohira et al., 2015; Lin et al., 2015; Le Goues et al., 2015; Just et al., 2014; Lamkanfi et al., 2013; D’Ambros et al., 2012; Zaman et al., 2012; Jalbert et al., 2011; Chen et al., 2021b; Selakovic and Pradel, 2017; Gunawi et al., 2014; Ho et al., 2023; Xu et al., 2023; Wang et al., 2023a; Zhao et al., 2023b; Wang et al., 2023c, 2022b; Shen et al., 2021; Wang et al., 2021a; Ren et al., 2020; Franco et al., 2017; Wang et al., 2017; Tan et al., 2018) 57
Version Control Systems (Timperley et al., 2024; Tebib et al., 2024; Yu et al., 2024; Pramod et al., 2024; Wang et al., 2024; Song et al., 2023; Zhong et al., 2023; Lee et al., 2024; An et al., 2023; Avula et al., 2023; Mahbub et al., 2023; Morovati et al., 2023; Chen et al., 2023b; Jiang et al., 2023; Applis and Panichella, 2023; Csuvik and Vidács, 2022; Kim et al., 2022; Zhang et al., 2022b; Bui et al., 2022; Song et al., 2022; Paltenghi and Pradel, 2022; Akimova et al., 2021; Gyimesi et al., 2021; Yuan et al., 2021; Zheng et al., 2021; Nikitopoulos et al., 2021; Kamienski et al., 2021; Karampatsis and Sutton, 2020; Widyasari et al., 2020; Wang et al., 2020; Benton et al., 2019; Radu and Nadi, 2019; Xu et al., 2019; Liu et al., 2019; Just et al., 2014; de Santana et al., 2024; Wang et al., 2023b; Azad et al., 2023; Eghbali and Pradel, 2020; Gu et al., 2019) 40
Vulnerability Databases (Ruan et al., 2024; Ni et al., 2024; Jiang et al., 2024; Wang et al., 2024; da Ponte et al., 2023; Pereira et al., 2022; Nikitopoulos et al., 2021; Fan et al., 2020; Lin et al., 2019; Gkortzis et al., 2018; Ku et al., 2007; Ye et al., 2023a; Cifuentes et al., 2009) 13
Other Pre-Existing Datasets (Kim and Hong, 2023; Bui et al., 2022; Kamienski et al., 2021; Sánchez et al., 2020; Zhang et al., 2020; Ferenc et al., 2018; Ye et al., 2023a) 7
Developer Q&A Forums (Zhang et al., 2024a; Zhao et al., 2023a; Morovati et al., 2023; Chen et al., 2022; Bagherzadeh et al., 2020; Zhang et al., 2018; Chen et al., 2021b; de Santana et al., 2024; Wang et al., 2022b) 9
Coding Assignments (Liu et al., 2024a; Tian et al., 2024; Wu et al., 2024; Tan et al., 2017; Le Goues et al., 2015; Lin et al., 2017) 6
CI/CD (Silva et al., 2024; Saavedra et al., 2024; Murphy et al., 2020; Tomassi et al., 2019; Madeiral et al., 2019; Madeyski and Kawalerowicz, 2017) 6
Defect Synthesis (Cai et al., 2024; Liang et al., 2023; Zhang et al., 2022a; Ngo et al., 2021; Bures et al., 2020; Tian et al., 2024) 6
@@ 4.1.1. Issue Reports
Several datasets collect defects directly from GitHub issues. These include gDefects4DL (Liang et al., 2022), HasBugs (Applis and Panichella, 2023), Andror2 (Wendland et al., 2021), and Bugs4Q (Zhao et al., 2023a). In addition to GitHub, Apache Jira serves as another popular bug-tracking system, specifically for the large-scale software projects hosted by the Apache Software Foundation. Datasets like Defects4J (Just et al., 2014), ApacheJIT (Keshavarz and Nagappan, 2022), Bugs.jar (Saha et al., 2018), and Ohira et al. (2015) source defects from Apache Jira, where developers log comprehensive bug reports, including detailed reproduction steps, resolutions, and in many cases, links to the relevant commits that introduce and fix the bugs.
These datasets leverage the wealth of defect information available in issue reports, providing robust resources for research on bug prediction, fixing strategies, and software maintenance practices. However, relying exclusively on issue reports also introduces certain limitations, such as insufficient source-code context and difficulties in defect reproduction.
@@ 4.1.2. Version Control Systems
The commit histories of version control systems offer a detailed view of software evolution, providing a rich source for collecting software defects. By analyzing commit messages and code changes, researchers can identify bug-fixing commits and their corresponding buggy versions. Several datasets leverage Git commit histories as their primary source of defects. For instance, BugsPHP (Pramod et al., 2024) mines commit pairs from GitHub to capture bug fixes in PHP projects, while Defectors (Mahbub et al., 2023) focuses on Python, extracting bug-inducing commits through a combination of rule-based and keyword-based approaches. FixJS (Csuvik and Vidács, 2022) sources bug-fixing in JavaScript commits, grouping them into function-level changes, and Minecraft (Avula et al., 2023) offers a dataset with precise bug location information mined from Git commit histories. Similarly, D2A (Zheng et al., 2021) conducts differential analysis on commit histories, using static analysis tools to identify bug-fixing commits.
Besides GitHub commits, other datasets gather commits from GitLab, such as BugsC++ (An et al., 2023), or utilize the GitHub archive for large-scale multilingual dataset construction, like MBF (Zhong et al., 2023). These varied sources ensure a broad and diverse collection of bug data across different platforms and programming languages. Collecting defects from version control systems allows dataset creators to capture real-world bugs in a systematic and structured manner, but this approach may overlook defects that are not fixed through a single commit and often require additional context to fully understand the defect.

### v2

@@ 3.2.5. Other Types of Defects
3.3. Programming Languages of Defective Software
Another way of categorizing datasets is by the programming language in which the defective code is written. 134/151 of our surveyed defect datasets focus on one or multiple programming languages. These languages include Java, Python, C/C++, JavaScript, Go, and others. Table 5 presents the distribution of programming languages across software defect datasets. Java stands out as the most represented language, with a total of 57 (57/151) datasets, followed by C/C++ (52 datasets, 52/151) and Python (39 datasets, 39/151). The dominance of these languages is likely due to their popularity in software development and due to the focus of the research community on these languages. Other languages, such as F#, Groovy, Haskell, and Julia, have fewer datasets available.
The distribution of languages in defect datasets has evolved over time. Figure 5 shows the trend for the top five programming languages: Java, C/C++, Python, JavaScript, and Go. These five languages also represent the top five most used languages on GitHub (Beuke, 2025), suggesting that dataset creation reflects real-world development trends. Java datasets have increased steadily over the past decade, while Python has gained significant popularity since 2017. JavaScript and Go have experienced peaks within the past five years, and C/C++ datasets have appeared consistently over time. It is important to note the surge in software defect datasets since 2023. While the general increase in publication volume contributes to this rise, another factor is the growing interest in domain- or type-specific defects. The surge is particularly pronounced for C/C++ datasets, with 14 papers in 2023 (a 250% increase from 2022), driven by growing interest in performance bugs (DBLP:journals/tse/ZhaoXBCL23; DBLP:conf/msr/AzadIHR23), security vulnerabilities (DBLP:conf/raid/0001DACW23; DBLP:conf/uss/XuLDDLWPM23), and bugs in emerging domains such as machine learning frameworks (DBLP:conf/icse/GuanXLLB23; DBLP:conf/icsm/HoMI0SKNR23), compilers (DBLP:conf/wcre/WangZRLJ23; DBLP:journals/infsof/WangBWYGS23), and quantum computing (DBLP:conf/qsw/ZhaoWLLZ23). The trend continues in 2024 with 13 C/C++ datasets, suggesting sustained research momentum in infrastructure-level defect analysis. Another contributing factor is the rising demand for large-scale datasets to train and evaluate machine learning and large language models, which has spurred automated bug mining and benchmark construction studies in 2023 (DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/kbse/AnKCYY23; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/aina/PonteRM23a; DBLP:conf/msr/MahbubSR23; DBLP:conf/raid/0001DACW23; DBLP:conf/icst/KimH23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/kbse/YeCG23). This trend will be further discussed in Section 6.
Figure 6. Programming languages of defects in each domain.
Figure 7. Types of defects for each programming language.
@@ 3.3. Programming Languages of Defective Software
Figure 7 shows the distribution of the most popular programming languages across different application domains. Python and C/C++ dominate machine learning-related defect datasets, and Java and Kotlin cover the major datasets for mobile applications. Compilers and runtime engines have a variety of programming languages, including C/C++, Rust, Python, JavaScript, and Go. Most blockchain-related defects are from Solidity smart contracts, except for one dataset that collects defects from blockchain clients (DBLP:conf/msr/KimKL22), which are implemented in Java, Go, C#, and Rust. In other domains, such as autonomous systems, quantum computing, scientific computing, and computational notebooks, C/C++ and Python are commonly used, whereas datasets for cloud computing, databases, and the internet of things involve Java, Go, C/C++, Python, and JavaScript.
@@
Defect types are not equally distributed across programming languages. Figure 7 shows the distribution of software defect datasets across different programming languages and defect types. Java, C/C++, Python, JavaScript, and Go are common across all defect types, while C#, Rust, Solidity, and PHP cover functional, security, and performance defects only. Kotlin and Ruby have the least coverage, with Kotlin covering only functional and security defects, and Ruby covering only functional and performance defects.
RQ1 Summary Over 60% of the datasets target general-purpose software, while the rest cover eleven specific domains including machine learning, compilers & runtime engines, mobile applications, and blockchain. Machine learning is the most popular specific domain with 14 datasets. Software defect datasets cover functional, security, performance, and concurrency defects. Functional bugs dominate (100/151), followed by security vulnerabilities (48/151), performance issues (33/151), and concurrency defects (11/151). Security and performance datasets are more common in domain-specific contexts such as compilers & runtime engines or blockchain. Java, C/C++, and Python appear in over 75% of all datasets, with JavaScript and Go showing increasing presence recently. Python and C/C++ predominate in machine learning and compilers & runtime engines, while Java and Kotlin dominate mobile applications. Language trends partially align with defect types: Java, C/C++, Python, JavaScript, and Go span all defect types, while other languages associate with specific types.
4. Dataset Construction (RQ2)
@@ 4.1. Source of Defects
The source of a defect refers to the origin or platform from which the defect data is collected. As listed in Table 6, there are diverse sources for dataset creators to mine defects from, such as issue reports, version control systems, vulnerability databases, developer Q&A forums, and others. Most of these sources are closely tied to the software development process, as illustrated in Figure 8, especially in software implementation, testing, deployment, and maintenance stages. Other sources, such as coding assignments and defect synthesis, offer a more controlled environment for collecting or creating defects. Among these sources, issue reports and version control systems are the most popular, accounting for 66 (66/151) and 50 (50/151) software defect datasets, respectively. Coding assignments, CI/CD, and defect synthesis are less frequently used, with no more than 8 (8/151) datasets utilizing each. The rest of this section provides a detailed discussion of the primary sources of defects in software defect datasets, highlighting their unique characteristics.
@@
Issue Reports (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/edcc/AndradeLV24; DBLP:journals/tosem/LiuMC24; DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:conf/icse/CuiD0WSZWYXH0024; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/issta/XuG024; DBLP:conf/ease/WaseemDA0M24; DBLP:conf/kbse/DasAM24; DBLP:conf/qrs/YuW24; DBLP:conf/iclr/JimenezYWYPPN24; DBLP:journals/jss/ZhaoMLZ23; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/icse/BhuiyanPVPS23; DBLP:conf/icse/GuanXLLB23; DBLP:conf/issta/XiongX0SWWP0023; DBLP:journals/softx/PachoulyAK22; DBLP:conf/icse/LiangLSSFD22; DBLP:conf/msr/KeshavarzN22; DBLP:conf/issre/LiuZDHDM022; DBLP:conf/msr/WendlandSMMHMRF21; DBLP:conf/msr/KimKL21; DBLP:conf/icse/PaulTB21a; DBLP:conf/issre/Du0MZ21; DBLP:conf/icse/Makhshari021; DBLP:journals/jss/FerencGGTG20; DBLP:journals/jss/WuZCWM20; DBLP:conf/icse/GarciaF0AXC20; DBLP:conf/promise/VieiraSRG19; DBLP:conf/msr/RiganelliMMM19; DBLP:conf/msr/PontaPSBD19; DBLP:conf/msr/SahaLLYP18; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/msr/OhiraKYYMLFHIM15; DBLP:conf/kbse/LinMZCZ15; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/issta/JustJE14; DBLP:conf/msr/LamkanfiPD13; DBLP:journals/ese/DAmbrosLR12; DBLP:conf/msr/ZamanAH12; DBLP:conf/hotpar/JalbertPPS11; DBLP:conf/icse/ChenYLCLWL21; DBLP:conf/se/SelakovicP17; DBLP:conf/cloud/GunawiHLPDAELLM14; DBLP:conf/icsm/HoMI0SKNR23; DBLP:conf/uss/XuLDDLWPM23; DBLP:journals/infsof/WangBWYGS23; DBLP:conf/qsw/ZhaoWLLZ23; DBLP:conf/wcre/WangZRLJ23; DBLP:conf/kbse/WangWCCY22; DBLP:conf/sigsoft/ShenM0TCC21; DBLP:conf/sigsoft/WangLX0S21; DBLP:conf/iwpc/RenL020; DBLP:conf/kbse/FrancoGR17; DBLP:conf/kbse/WangDGGQYW17; DBLP:conf/icse/TanDGR18; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/OliverDAH24; DBLP:conf/sigsoft/GaoDQGW0HZW18; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/access/AntalVKMHF24; DBLP:journals/ese/TambonNAKA24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tse/ZhaoXBCL23) 66
Version Control Systems (DBLP:journals/ese/TimperleyHSDW24; DBLP:conf/pst/TebibAAG24; DBLP:conf/issta/YuXZ0LS24; DBLP:conf/msr/PramodSTSW24; DBLP:conf/icse/WangHGWC024; DBLP:conf/kbse/SongWCCLLWP23; DBLP:conf/kbse/ZhongLZGWGL23; DBLP:conf/icst/LeeKYY24; DBLP:conf/kbse/AnKCYY23; DBLP:conf/kbse/AvulaVM23; DBLP:conf/msr/MahbubSR23; DBLP:journals/ese/MorovatiNKJ23; DBLP:conf/raid/0001DACW23; DBLP:journals/tse/JiangLLZCNZHBZ23; DBLP:conf/msr/ApplisP23; DBLP:conf/msr/CsuvikV22; DBLP:conf/msr/KimKL22; DBLP:conf/wcre/ZhangYYCYZ22; DBLP:conf/msr/BuiSF22; DBLP:conf/issta/Song0NW0DM22; DBLP:journals/pacmpl/PaltenghiP22; DBLP:conf/apsec/AkimovaBDKKMM21; DBLP:journals/stvr/GyimesiVSMBFM21; DBLP:conf/cgo/YuanLLLLX21; DBLP:conf/icse/ZhengPLBEYLMS21; DBLP:conf/sigsoft/NikitopoulosDLM21; DBLP:conf/msr/KamienskiPBH21; DBLP:conf/msr/KarampatsisS20; DBLP:conf/sigsoft/WidyasariSLQPTT20; DBLP:conf/msr/WangBJS20; DBLP:conf/icse/BentonGZ19; DBLP:conf/msr/RaduN19; DBLP:conf/dsa/XuYWA19; DBLP:journals/ese/LiuWWXCWYZ19; DBLP:conf/issta/JustJE14; DBLP:journals/tosem/SantanaNAA24; DBLP:journals/jss/WangCHZBZ23; DBLP:conf/msr/AzadIHR23; DBLP:conf/kbse/EghbaliP20; DBLP:conf/compsac/GuWL0019; DBLP:conf/icse/HumbatovaJBR0T20; DBLP:conf/issta/WuJPLD0BS23; DBLP:conf/kbse/RomanoLK021; DBLP:conf/msr/AmannNNNM16; DBLP:conf/msr/BeyerGKRR24; DBLP:journals/access/AntalVKMHF24; DBLP:journals/pacmpl/DrososSAM024; DBLP:journals/tosem/TufanoWBPWP19; DBLP:journals/tse/ZhaoXBCL23; DBLP:conf/kbse/XuZL23) 50
Vulnerability Databases (DBLP:conf/kbse/RuanLZL24; DBLP:conf/msr/NiSYZW24; DBLP:journals/tosem/JiangJWMLZ24; DBLP:conf/icse/WangHGWC024; DBLP:conf/aina/PonteRM23a; DBLP:conf/prdc/PereiraAV22; DBLP:conf/sigsoft/NikitopoulosDLM21; DBLP:conf/msr/FanL0N20; DBLP:conf/icics/LinXZX19; DBLP:conf/msr/GkortzisMS18; DBLP:conf/kbse/KuHCL07; DBLP:conf/kbse/YeCG23; BegBunch; DBLP:conf/icse/HuZLYH024; DBLP:conf/issta/WuJPLD0BS23) 15
Other Pre-Existing Datasets (DBLP:conf/icst/KimH23; DBLP:conf/msr/BuiSF22; DBLP:conf/msr/KamienskiPBH21; DBLP:journals/access/SanchezDMS20; DBLP:conf/icsm/ZhangXL20; DBLP:conf/promise/FerencTLSG18; DBLP:conf/kbse/YeCG23; DBLP:conf/msr/AmannNNNM16; DBLP:conf/msr/OliverDAH24; DBLP:conf/sigsoft/GaoDQGW0HZW18) 10
Developer Q&A Forums (DBLP:journals/tosem/ZhangCWCLMMHL24; DBLP:journals/jss/ZhaoMLZ23; DBLP:journals/ese/MorovatiNKJ23; DBLP:journals/tse/ChenXLGLC22; DBLP:journals/pacmpl/0001FSK20; DBLP:conf/issta/ZhangCCXZ18; DBLP:conf/icse/ChenYLCLWL21; DBLP:journals/tosem/SantanaNAA24; DBLP:conf/kbse/WangWCCY22; DBLP:conf/icse/HumbatovaJBR0T20) 10
CI/CD (DBLP:conf/msr/SilvaSM24; DBLP:conf/icse/SaavedraSM24; DBLP:conf/hotsos/MurphyBSR20; DBLP:conf/icse/DmeiriTWBLDVR19; DBLP:conf/wcre/DelfimUMM19; DBLP:conf/msr/MadeyskiK17; DBLP:conf/sigsoft/Zhang0C0Z19; DBLP:conf/wcre/ReyesGSBM24) 8
Defect Synthesis (DBLP:conf/acl/TianYQCLPWHL0024; DBLP:journals/jss/CaiLZZS24; DBLP:conf/asiaccs/LiangYDMHZ23; DBLP:conf/uss/ZhangP0W22; DBLP:conf/splc/NgoNNV21; DBLP:conf/icst/BuresHA20; DBLP:conf/sp/Dolan-GavittHKL16; DBLP:conf/icse-apr/YadavW24) 8
Coding Assignments (DBLP:conf/msr/LiuHLZCSHM24; DBLP:conf/acl/TianYQCLPWHL0024; DBLP:conf/sigsoft/WuLZ024; DBLP:conf/icse/TanYYMR17; DBLP:journals/tse/GouesHSBDFW15; DBLP:conf/oopsla/LinKCS17; DBLP:conf/fie/SandersWA24) 7
@@ 4.1.1. Issue Reports
Several datasets collect defects directly from GitHub issues. These include gDefects4DL (DBLP:conf/icse/LiangLSSFD22), HasBugs (DBLP:conf/msr/ApplisP23), Andror2 (DBLP:conf/msr/WendlandSMMHMRF21), and Bugs4Q (DBLP:journals/jss/ZhaoMLZ23). In addition to GitHub, Apache Jira serves as another popular bug-tracking system, specifically for the large-scale software projects hosted by the Apache Software Foundation. Datasets like Defects4J (DBLP:conf/issta/JustJE14), ApacheJIT (DBLP:conf/msr/KeshavarzN22), Bugs.jar (DBLP:conf/msr/SahaLLYP18), and DBLP:conf/msr/OhiraKYYMLFHIM15 source defects from Apache Jira, where developers log comprehensive bug reports, including detailed reproduction steps, resolutions, and in many cases, links to the relevant commits that introduce and fix the bugs. Bugzilla is another prominent bug-tracking system that has been leveraged as a defect source by several software defect datasets, including the Eclipse and Mozilla Defect Tracking Dataset (DBLP:conf/msr/LamkanfiPD13), BugHub (DBLP:conf/edcc/AndradeLV24), and CISB (DBLP:conf/uss/XuLDDLWPM23).
These datasets leverage the wealth of defect information available in issue reports, providing robust resources for research on bug prediction, fixing strategies, and software maintenance practices. However, relying exclusively on issue reports introduces several limitations. First, many issue reports lack explicit links to the commits or code regions that caused or fixed the defect, making it difficult to establish accurate traceability between natural-language descriptions and source-code changes (Herzig et al., 2013; Zhou et al., 2012). Second, issue reports are often incomplete or poorly described, missing key environmental or configuration information, which complicates defect reproduction and automated analysis (Bettenburg et al., 2008; Zimmermann et al., 2010). Consequently, datasets built from issue reports typically require extensive manual curation to ensure correctness and reproducibility.
@@ 4.1.2. Version Control Systems
The commit histories of version control systems offer a detailed view of software evolution, providing a rich source for collecting software defects. By analyzing commit messages and code changes, researchers can identify bug-fixing commits and their corresponding buggy versions. Several datasets leverage Git commit histories as their primary source of defects. For instance, BugsPHP (DBLP:conf/msr/PramodSTSW24) mines commit pairs from GitHub to capture bug fixes in PHP projects, while Defectors (DBLP:conf/msr/MahbubSR23) focuses on Python, extracting bug-inducing commits through a combination of rule-based and keyword-based approaches. FixJS (DBLP:conf/msr/CsuvikV22) sources bug-fixing in JavaScript commits, grouping them into function-level changes, and Minecraft (DBLP:conf/kbse/AvulaVM23) offers a dataset with precise bug location information mined from Git commit histories. Similarly, D2A (DBLP:conf/icse/ZhengPLBEYLMS21) conducts differential analysis on commit histories, using static analysis tools to identify bug-fixing commits.
Besides GitHub commits, other datasets gather commits from GitLab, such as BugsC++ (DBLP:conf/kbse/AnKCYY23), or utilize the GitHub archive for large-scale multilingual dataset construction, like MBF (DBLP:conf/kbse/ZhongLZGWGL23). These varied sources ensure a broad and diverse collection of bug data across different platforms and programming languages. Collecting defects from version control systems allows dataset creators to capture real-world bugs systematically, but this approach faces challenges. For example, identifying which commits actually introduce or fix defects is difficult, as approaches such as SZZ (Sliwerski et al., 2005, 2005; Kim et al., 2006) often misclassify refactorings or multi-commit fixes (Neto et al., 2018). Additionally, commit messages vary widely in quality, making keyword-based mining unreliable (Tian et al., 2012) and often requiring additional context to fully understand the defect.